# A Metamorphic Testing Perspective on Knowledge Distillation for Language Models of Code: Does the Student Deeply Mimic the Teacher?

### v2

@@ 1 Introduction
We investigate the student model’s ability to mimic the teacher by performing simple yet effective adversarial attacks Yang et al. [2022], Zhang et al. [2020a], Zeng et al. [2022] that rename identifiers based on the contextual semantics of each code snippet. The empirical results reveal that although the student models achieve performance comparable to the teacher before adversarial perturbations, their performance degrades substantially under adversarial attacks, with up to 285% greater performance loss than the teacher model. This finding indicates that the student model does not deeply mimic the teacher’s predictive behavior and internal representations, highlighting the limitations of the traditional accuracy-based evaluation, which often overlooks more profound behavioral discrepancies between teacher and student models. It also underscores the need for a nuanced evaluation framework that goes beyond accuracy to assess how deeply a student model mimics the predictive behavior and internal representations of its teacher, providing a more comprehensive understanding of model alignment under knowledge distillation.
@@ 2.3 Metamorphic Testing
Metamorphic testing (MT) is a property-based testing paradigm widely adopted for validating software systems in the absence of a test oracle or when one is difficult to define, a challenge commonly referred to as the oracle problem Chen et al. [2018], Zhang et al. [2014]. Rather than verifying the correctness of individual outputs in isolation, MT addresses this issue by evaluating whether the program under test exhibits expected relationships between outputs across a set of well-defined metamorphic relations (MRs). An MR specifies a necessary property that captures the expected relationship between multiple inputs and their corresponding outputs. These relations are derived from the problem domain’s inherent semantics and define how the output should change (or remain consistent) when the input is systematically modified. This paradigm is beneficial for testing complex systems such as machine learning models, scientific computing software, or simulation tools, where correctness is difficult to assert for individual outputs but consistency across related cases can be meaningfully assessed Yang et al. [2025], Xie et al. [2011], Chen et al. [2020], Xiao et al. [2022], Segura et al. [2018].
@@ 3.1.2 Datasets and Subjects
Clone detection: (CD) This task aims to determine whether two given code snippets are syntactically or semantically equivalent. To streamline the fine-tuning process, we adopt the filtered BigCloneBench dataset Svajlenko et al. [2014] provided by Wang et al. Wang et al. [2020]. The final dataset includes 90,102 training examples, along with 4,000 examples each for validation and testing. Vulnerability prediction: (VP) This task aims to determine whether a given code snippet contains a vulnerability. We conduct our experiments using the Devign [^5] https://sites.google.com/view/devign dataset, introduced by Zhou et al. Zhou et al. [2019]. The dataset comprises 27,318 functions sourced from two widely used open-source C libraries, FFmpeg and Qemu. Each function is manually annotated to indicate whether it contains vulnerabilities. We use the CodeXGLUE benchmark splits Lu et al. [2021], including training, development, and test sets with 27,318, 2,732, and 2,732 examples, respectively.
CodeBERT is a pre-trained language model that shares the same architecture and a similar pre-training strategy as RoBERTa Liu [2019]. It consists of twelve Transformer layers, each with 12-headed multi-head attention mechanisms and a hidden size of 768. CodeBERT is pre-trained using two objectives: masked language modeling, which predicts masked tokens, and replaced token detection, which identifies substituted tokens in the input. The model is trained on bimodal data comprising natural language–code pairs across six programming languages from the CodeSearchNet dataset Husain et al. [2019]. GraphCodeBERT shares CodeBERT’s architecture but incorporates data-flow information from source code during pre-training. Additionally, it introduces a pre-training task that predicts whether a data flow edge exists between two nodes.
@@ 3.1.3 Adversarial Attack Approaches
In this study, adversarial attacks are employed to evaluate how deeply the student model mimics the teacher’s predictive behavior and internal representations during knowledge distillation. Adversarial attacks are typically categorized as black-box or white-box, depending on the attacker’s level of access to the model. Following the recommendation of Du et al. Du et al. [2023], we adopt black-box attacks, as they rely solely on model outputs and are widely used for evaluating language models of code Yang et al. [2022], Zhang et al. [2020a]. The perturbations introduced through these attacks are deliberately subtle to maintain the syntactic and semantic integrity of the source code, as discussed in Section 2.4, while inducing mispredictions. To this end, we employ three established black-box techniques: ALERT Yang et al. [2022], Metropolis Hastings Modifier (MHM) Zhang et al. [2020a], and WIR-Random Zeng et al. [2022], each of which performs identifier substitution to assess model robustness without altering code functionality. ALERT Yang et al. [2022] considers both operational and natural semantics, using greedy and genetic algorithms to select substitutions, with 30 candidates per identifier. MHM Zhang et al. [2020a] utilizes Metropolis-Hastings sampling for iterative identifier substitution, with a maximum of 100 iterations and 30 candidates per iteration. WIR-Random Zeng et al. [2022] ranks identifiers based on their impact on model predictions and replaces them sequentially with randomly chosen candidates. It is important to note that only instances correctly classified before perturbation are included, ensuring that the robustness evaluation is based exclusively on valid predictions and that any observed degradation in model performance is caused solely by adversarial perturbations rather than pre-existing misclassifications Yang et al. [2022], Zhang et al. [2020a].
@@ 4.1 Behavioral Fidelity as a Software Testing Problem
As a result, conventional software testing techniques such as regression testing, which depend on exact output matches, are often too rigid—minor differences in probabilistic outputs or generated sequences do not necessarily indicate functional errors. Mutation testing is also ill-suited, as neither model is inherently faulty. In contrast, metamorphic testing provides a principled approach to overcoming the oracle problem by defining expected relations (as detailed in Section 4.2) between the outputs of the teacher and student models, without requiring explicit ground-truth. These relations act as relational oracles, enabling us to verify whether the teacher and student models produce consistent outputs to the same inputs. Finally, metamorphic testing has been widely adopted for evaluating machine learning–based software systems and large language models Xie et al. [2020], Sun et al. [2024], Kwan et al. [2024], Zhang et al. [2020b], Wu et al. [2025], Yang et al. [2025], demonstrating its versatility and effectiveness in oracle-free evaluation scenarios. Building on this foundation, MetaCompress adopts a metamorphic testing–based approach to evaluate behavioral fidelity of student models, framing it as an oracle-free testing problem.

### v3

@@ 1 Introduction
We investigate the student model’s ability to mimic the teacher by performing simple yet effective adversarial attacks Yang et al. [2022], Zhang et al. [2020a], Zeng et al. [2022], Tian et al. [2023]. These attacks involve renaming identifiers and applying semantic-preserving code transformations informed by the contextual semantics of each code snippet. The empirical results reveal that although the student models achieve performance comparable to the teacher before adversarial perturbations, their performance degrades substantially under adversarial attacks, with up to 285% greater performance loss than the teacher model. This finding indicates that the student model does not deeply mimic the teacher’s predictive behavior and internal representations, highlighting the limitations of the traditional accuracy-based evaluation, which often overlooks more profound behavioral discrepancies between teacher and student models. It also underscores the need for a nuanced evaluation framework that goes beyond accuracy to assess how deeply a student model mimics the predictive behavior and internal representations of its teacher, providing a more comprehensive understanding of model alignment under knowledge distillation.
@@ 2.3 Metamorphic Testing
Metamorphic testing (MT) is a property-based testing paradigm widely adopted for validating software systems in the absence of a test oracle or when one is difficult to define, a challenge commonly referred to as the oracle problem Chen et al. [2018], Zhang et al. [2014]. Rather than verifying the correctness of individual outputs in isolation, MT addresses this issue by evaluating whether the program under test exhibits expected relationships between outputs across a set of well-defined metamorphic relations (MRs). An MR specifies a necessary property that captures the expected relationship between multiple inputs and their corresponding outputs. These relations are derived from the problem domain’s inherent semantics and define how the output should change (or remain consistent) when the input is systematically modified. This paradigm is beneficial for testing complex systems such as machine learning models, scientific computing software, or simulation tools, where correctness is difficult to assert for individual outputs but consistency across related cases can be meaningfully assessed Yang et al. [2025a], Xie et al. [2011], Chen et al. [2020], Xiao et al. [2022], Segura et al. [2018].
@@ 3.1.2 Datasets and Subjects
Clone detection (CD): This task aims to determine whether two given code snippets are syntactically or semantically equivalent. To streamline the fine-tuning process, we adopt the filtered BigCloneBench dataset Svajlenko et al. [2014] provided by Wang et al. Wang et al. [2020]. The final dataset includes 90,102 training examples, along with 4,000 examples each for validation and testing.
Vulnerability prediction (VP): This task aims to determine whether a given code snippet contains a vulnerability. We conduct our experiments using the Devign [^5] https://sites.google.com/view/devign dataset, introduced by Zhou et al. Zhou et al. [2019]. The dataset comprises 27,318 functions sourced from two widely used open-source C libraries, FFmpeg and Qemu. Each function is manually annotated to indicate whether it contains vulnerabilities. We use the CodeXGLUE benchmark splits Lu et al. [2021], including training, development, and test sets with 27,318, 2,732, and 2,732 examples, respectively.
CodeBERT is a pre-trained language model that shares the same architecture and a similar pre-training strategy as RoBERTa Liu [2019]. It consists of twelve Transformer layers, each with 12-headed multi-head attention mechanisms and a hidden size of 768. CodeBERT is pre-trained using two objectives: masked language modeling, which predicts masked tokens, and replaced token detection, which identifies substituted tokens in the input. The model is trained on bimodal data comprising natural language–code pairs across six programming languages from the CodeSearchNet dataset Husain et al. [2019].
GraphCodeBERT shares CodeBERT’s architecture but incorporates data-flow information from source code during pre-training. Additionally, it introduces a pre-training task that predicts whether a data flow edge exists between two nodes.
@@ 3.1.3 Adversarial Attack Approaches
In this study, adversarial attacks are employed to evaluate how deeply the student model mimics the teacher’s predictive behavior and internal representations during knowledge distillation. Adversarial attacks are typically categorized as black-box or white-box, depending on the attacker’s level of access to the model. Following the recommendation of Du et al. Du et al. [2023], we adopt black-box attacks, as they rely solely on model outputs and are widely used for evaluating language models of code Yang et al. [2022], Zhang et al. [2020a]. The perturbations introduced through these attacks are deliberately subtle to maintain the syntactic and semantic integrity of the source code, as discussed in Section 2.4, while inducing mispredictions. To this end, we employ three established black-box techniques: ALERT Yang et al. [2022], Metropolis Hastings Modifier (MHM) Zhang et al. [2020a], and WIR-Random Zeng et al. [2022], each of which performs identifier substitution to assess model robustness without altering code functionality. To further enhance robustness evaluation, we also employ CODA Tian et al. [2023], a black-box attack technique that combines identifier renaming with semantic-preserving code transformations. This enables assessment of model robustness under a wider range of functionally equivalent program variations.
ALERT considers both operational and natural semantics, using greedy and genetic algorithms to select substitutions, with 30 candidates per identifier. MHM utilizes Metropolis-Hastings sampling for iterative identifier substitution, with a maximum of 100 iterations and 30 candidates per iteration. WIR-Random ranks identifiers based on their impact on model predictions and replaces them sequentially with randomly chosen candidates. In contrast to these identifier-centric techniques, CODA utilizes code differences between a target input and reference inputs with divergent predictions to guide adversarial example generation. It combines identifier renaming with semantic-preserving structural transformations. By considering both the identifier level and structural differences, CODA generates functionally equivalent yet more diverse adversarial examples.
It is important to note that only instances correctly classified before perturbation are included, ensuring that the robustness evaluation is based exclusively on valid predictions and that any observed degradation in model performance is caused solely by adversarial perturbations rather than pre-existing misclassifications Yang et al. [2022], Zhang et al. [2020a].
@@ 3.2 Empirical Results
To further assess the robustness of teacher and student models to a broader class of semantic-preserving adversarial perturbations, we conduct additional experiments using the CODA attack. Figure 3 presents the Attack Success Rate values under CODA. The results align with the trends observed for ALERT, MHM, and WIR-Random attacks. Across both tasks and model architectures, student models trained via knowledge distillation generally achieve higher ASR values than their teacher counterparts, indicating reduced robustness to CODA-induced perturbations. This finding applies to both AVATAR and MORPH students, confirming that the previously observed robustness degradation is not restricted to identifier-level attacks.
Figure 3: Attack success rate metric values across CODA adversarial attack, models, tasks, and knowledge distillation methods.
@@ 4.1 Behavioral Fidelity as a Software Testing Problem
As a result, conventional software testing techniques such as regression testing, which depend on exact output matches, are often too rigid—minor differences in probabilistic outputs or generated sequences do not necessarily indicate functional errors. Mutation testing is also ill-suited, as neither model is inherently faulty. In contrast, metamorphic testing provides a principled approach to overcoming the oracle problem by defining expected relations (as detailed in Section 4.2) between the outputs of the teacher and student models, without requiring explicit ground-truth. These relations act as relational oracles, enabling us to verify whether the teacher and student models produce consistent outputs to the same inputs. Finally, metamorphic testing has been widely adopted for evaluating machine learning–based software systems and large language models Xie et al. [2020], Sun et al. [2024], Kwan et al. [2024], Zhang et al. [2020b], Wu et al. [2025], Yang et al. [2025a], demonstrating its versatility and effectiveness in oracle-free evaluation scenarios. Building on this foundation, MetaCompress adopts a metamorphic testing–based approach to evaluate behavioral fidelity of student models, framing it as an oracle-free testing problem.
@@ 5.1 Behavioral Fidelity and Knowledge Gap in Knowledge Distillation
5.2 Practical Consideration and Deployment Cost
From a practical standpoint, evaluating the behavioral fidelity does not impose significant overhead. MetaCompress is designed exclusively as an evaluation framework to determine whether a distilled student model deeply mimics its teacher prior to deployment. In standard knowledge distillation workflows, both the teacher and student models are evaluated on held-out test data to assess predictive performance, which requires forward passes through both models. MetaCompress leverages the probability distributions and related outputs generated during this routine evaluation as byproducts, without requiring additional model executions. Therefore, even for teams that frequently conduct model distillation, MetaCompress introduces only negligible computational overhead beyond lightweight post-processing. This approach is both practical and scalable for diagnosing behavioral fidelity gaps and ensuring that compressed student models are robust and reliable for real-world deployment.
5.3 Actionable Insights for Practitioners and Researchers
Although MetaCompress reports violation rates as diagnostic measures of behavioral fidelity divergence, these values should not be interpreted in isolation. Instead, they serve as actionable indicators that inform decision-making throughout the model development and deployment pipeline.
5.3.1 Insights for Practitioners
  1. 1.
    The appropriate response to an observed metamorphic relation (MR) violation rate depends on the intended deployment context and the acceptable level of risk. For example, a student model with a low violation rate (e.g., below 10%) may be considered sufficiently stable for use in non-critical applications, such as software clone detection. In contrast, for security-critical tasks (e.g., vulnerability detection) or adversarially exposed environments, practitioners should exercise greater caution and avoid deploying highly compact models.
  2. 2.
    The acceptable MR violation rate is context-dependent and varies according to the risk tolerance of the specific application. It is recommended that practitioners integrate violation-rate checks into continuous integration and deployment pipelines to serve as a gating criterion for model release. Models exhibiting low violation rates may be suitable for deployment in non-critical applications. In contrast, models designed for security-sensitive or adversarial environments should adhere to more stringent thresholds or be supplemented with additional safeguards, such as adversarial defenses or post-deployment monitoring, prior to release.
  3. 3.
    When violation rates indicate substantial behavioral divergence between teacher and student models within a specific application context, practitioners may consider revisiting the selected student architecture or knowledge distillation strategy. Adopting alternative student configurations, improved supervision mechanisms, or multi-teacher distillation approaches may mitigate metamorphic relation violations and reduce the knowledge gap, thereby enhancing behavioral fidelity prior to deployment.
5.3.2 Insights for Researchers
  1. 1.
    From a research perspective, MR violation rates provide a quantitative signal for diagnosing knowledge transfer quality beyond conventional accuracy metrics. Since student models exhibit lower adversarial robustness than their teacher counterparts, the soft knowledge transferred from the teacher to the student does not adequately reflect the teacher’s decision boundary; hence, the researcher should focus on improving the quality of the transferred knowledge.
  2. 2.
    Yang et al. Yang et al. [2025b] demonstrated that Mixture of Experts (MoE) frameworks enhance software vulnerability detection relative to single-model approaches. Building on this finding, subsequent research could investigate integrating MoE architectures into knowledge distillation to extract complementary knowledge from multiple specialized teachers, rather than relying on a single generalist model.
  3. 3.
    In this study, the selected knowledge distillation approaches focus on aligning the final-layer outputs of the teacher and student models as the primary training objective. While this design choice is widely adopted, future research may explore incorporating intermediate-layer alignment as an additional or complementary objective to further reduce the knowledge gap between teacher and student models and potentially improve behavioral fidelity.
In summary, the insights above suggest that behavioral fidelity divergence and metamorphic relation violation rates should be interpreted as decision-support metrics rather than pass–fail criteria. By making decision-boundary instability observable and measurable, MetaCompress enables developers to make informed follow-up decisions—such as pursuing architectural refinements or accepting a model based on application-specific risk tolerance—and encourages researchers in designing more robust and reliable knowledge distillation techniques.